OpenStudy (anonymous):

what is radical 8 times radical 3

jimthompson5910 (jim_thompson5910):

sqrt(8)*sqrt(3) = sqrt(8*3) = sqrt(24) So sqrt(8)*sqrt(3) = sqrt(24) ------------------------------------------ sqrt(24) = sqrt(4*6) = sqrt(4)*sqrt(6) = 2*sqrt(6) which means sqrt(24) = 2*sqrt(6) So in the end, sqrt(8)*sqrt(3) = 2*sqrt(6)

OpenStudy (anonymous):

@jim_thompson5910 thank you very much!:)

jimthompson5910 (jim_thompson5910):

you're welcome